OpenStudy (anonymous):

The equation (x + 6)^2 + (y + 4)^2 = 36 models the position and range of the source of a radio signal. Describe the position of the source and the range of the signals.

OpenStudy (binary3i):

it is a circle. range is the radius and source is the center. do you know how to bring them out of equation?

OpenStudy (anonymous):

No :/

OpenStudy (campbell_st):

the general form of a circle is (x -h)^2 + (y = k)^2 = r^2 the centre is (h, k) and radius is r in your question h = -6, k = -4 and r = 6 hopefully this helps...

OpenStudy (binary3i):

@campbell_st is correct.

OpenStudy (campbell_st):

oops (x -h)^2 + (y - k)^2 = r^2

OpenStudy (binary3i):

now he is more correct

OpenStudy (anonymous):

I got it;